### v3

# CommonPower: A Framework for Safe Data-Driven Smart Grid Control

Abstract
The growing complexity of power system management has led to an increased interest in reinforcement learning (RL). However, vanilla RL controllers cannot themselves ensure satisfaction of system constraints. Therefore, combining them with formally correct safeguarding mechanisms is an important aspect when studying RL for power system management. Integrating safeguarding into complex use cases requires tool support. To address this need, we introduce the Python tool CommonPower. CommonPower’s unique contribution lies in its symbolic modeling approach, which enables flexible, model-based safeguarding of RL controllers. Moreover, CommonPower offers a unified interface for single-agent RL, multi-agent RL, and optimal control, with seamless integration of different forecasting methods. This allows users to validate the effectiveness of safe RL controllers across a large variety of case studies and investigate the influence of specific aspects on overall performance. We demonstrate CommonPower’s versatility through a numerical case study that compares RL agents featuring different safeguards with a model predictive controller in the context of building energy management.
Index Terms:
Safe reinforcement learning, energy management, model predictive control, multi-agent systems, and forecast uncertainties.
I Introduction
I-A Motivation
Increasing adoption of intermittent renewable energy generation and complex power demand patterns, e.g., from electrifying heating and mobility, challenge power system control. Reinforcement learning (RL) has emerged as a promising method as it does not require explicit model knowledge and can automatically adapt to changing parameters. RL controllers have successfully been demonstrated for Volt-Var control, frequency control, economic dispatch, and smart home energy management [1]. Additionally, especially in the current transition period in which not every prosumer has sufficiently smart monitoring infrastructure, distributed control based on multi-agent reinforcement learning (MARL) provides a versatile data-driven approach largely using local information [2, 3, 4, 5].
RL controllers have demonstrated competitive performance compared to model predictive controllers (MPC) at a significantly lower computational cost [6]. Furthermore, there is evidence that RL controllers can learn the implicit patterns of exogenous inputs and thus outperform a nominal MPC in a setting with inaccurate forecasts [7]. For all its promise, integrating RL controllers into smart grids faces a major challenge since vanilla RL cannot guarantee the satisfaction of system constraints.
Safe RL is an active area of research with a large amount of existing literature [8, 9, 10, 11]. Three different safety levels can be distinguished: Encouraged constraint satisfaction, constraint satisfaction with high probability, and guaranteed constraint satisfaction [10]. Due to the critical nature of power systems, guaranteed constraint satisfaction is the required safety level for the real-world deployment of RL controllers. Some approaches require finding Lyapunov functions to prove stability [12] or control barrier functions to prove safety [13] – these are very hard to find for non-experts, especially when the system dynamics are complex.
Therefore, we focus on approaches that can readily establish safety guarantees based on a possibly significantly simplified system model [10, 14]. Such a model must be an abstraction of reality in that it produces more behaviors than the real system and can be computed via reachset-conformant identification [15, 16]. Recent work has shown that the specific type of safeguarding and the corresponding feedback given to the RL agent can influence the control performance, which motivates further research [14, 17].
A whole range of Python libraries have been published for modeling power systems and interfacing RL agents. However, these tools have hard-coded safety mechanisms that do not provide feedback to the RL controller. At the same time, sophisticated RL safeguarding strategies are an active area of research for specific applications but have not yet been included in a generic tool.
Our Python library CommonPower [^1] https://github.com/TUMcps/commonpower closes this gap and addresses the need for a versatile tool facilitating the exploration of safe RL controllers in a large variety of smart grid control use cases. By facilitating research on optimized safeguarding approaches, providing credible optimal control baselines, and modeling complex multi-agent setups, CommonPower lays the foundation for increased real-world adoption of RL for smart grid control.
I-B Contributions
The central aim of CommonPower is to pave the way toward comprehensive analysis of safe RL-based controllers in the smart grid context. This work introduces the required theoretical abstractions, describes the corresponding implementation in CommonPower, and reports a case study on the influence of different safeguarding approaches. Specifically, our main contributions are:
  • Formally correct model-based safeguarding: The presence of inaccurate forecasts must be explicitly considered in the safeguarding mechanisms. We provide a general formalization for model-based safeguarding robustified with approaches from reachability analysis.
  • Control on arbitrary levels of decentralization: A coalition-based formulation enables unified treatment of single-agent and multi-agent RL and the combination of RL-based and non-RL controllers in the same system.
  • Modular Python library CommonPower: CommonPower facilitates component modeling via symbolic representations. These are used to automatically extract system constraints for safeguarding and a built-in model predictive controller. Furthermore, the library implements a unified interface for single-agent and multi-agent RL algorithms via a gymnasium interface [18]. Due to the highly modular structure, custom controllers, safeguarding mechanisms, and forecasters can be integrated purely by extension.
  • Numerical benchmark of different safeguarding approaches: We utilize CommonPower to show how different safeguarding types affect the performance of an RL agent tasked with the optimal scheduling of a battery.
Existing libraries that provide interfaces to RL-based controllers do not provide the modularity that would facilitate straightforward investigations along the above-mentioned dimensions. The comparison in Tab. I shows that CommonPower is the only tool in the smart grid context that provides flexible component modeling, symbolic equations, decentralized control, and flexible prediction models.
CommonPower is extensible on various levels. It supports defining new models and controllers through well-documented interfaces and tutorials. In addition, to facilitate integration into ongoing projects, an import interface to the well-known library pandapower [19] is provided.
The type of setup that can unprecedentedly be modeled in CommonPower is illustrated in Fig. 1. The figure shows decentral control with heterogeneous controllers and safeguarding based on symbolic representations.
Figure 1: Illustration of a decentral control set-up featuring non-RL and RL-based controllers with modular safeguards.
I-C Organization
After discussing related work (Sec. II), we introduce the high-level problem formulation and establish some theoretical RL background (Sec. III). We describe how power systems and forecasters are modeled in CommonPower (Sec. IV) and present our unified approaches to control and safeguarding (Sec. V). Finally, we investigate an exemplary case study (Sec. VI) and conclude (Sec. VII).
II Related Work
While many tools exist for modeling and controlling power systems using RL agents, none implement a systematic and versatile approach to safeguarding RL. Therefore, we first provide an overview of research investigating such strategies in specific applications. Subsequently, we present existing tools featuring single-agent and multi-agent RL for smart grid control.
II-A Control Applications Utilizing Safe Single-Agent RL
A possible strategy to safeguard a central RL controller is to apply an action mask that only lets the agent choose safe actions [20, 21, 22]. While this approach is well-established for discrete action spaces, it has, to the best of our knowledge, not been investigated for continuous action spaces in the smart grid context.
The most common strategy for continuous action spaces is to let the agent freely choose an action and adjust this proposed action before applying it to the system. In the area of voltage regulation, the studies in [23, 24] formulate a projection based on a quadratic optimal program with linearized voltage constraints. The work in [25] safeguards an economic dispatch controller with an efficient linear projection, while the studies in [6, 26] use a quadratic program, [26] additionally including a constraint that guarantees demand satisfaction during islanding mode.
However, the study in [27] argues that such projections might have unfavorable implications on the learning performance of the RL agent, which demands further investigation. A step in this direction is the incorporation of a differentiable projection layer within the policy network of the RL agent [28, 29]. In a similar fashion, the approach in [30] adjusts the output of the policy network in two steps. The first completes the partial network output to satisfy equality constraints, while the second computes a correction to additionally satisfy inequality constraints. In contrast, the work in [31] develops a differentiable function that maps the action of the RL agent to a previously computed robust control invariant set.
II-B Control Applications Utilizing Safe Multi-Agent RL
In multi-agent RL for power system control, we need to differentiate between safeguarding constraints that depend on the behavior of multiple agents, which we refer to as global constraints, and constraints that only apply to an individual agent, referred to as local constraints.
Local constraints, such as the charge limits of independent energy storage systems, can be addressed using the approaches presented in the previous subsection but are commonly neglected in the multi-agent literature. Some works assume underlying backup controllers such as a battery management system that will override unsafe decisions of individual agents [32, 33]. The study in [17] employs a quadratic program to guarantee local constraint satisfaction.
Global constraints whose satisfaction depends on the decisions of all agents are generally modeled as soft constraints: If these constraints are violated, all agents receive a negative reward encouraging them to reduce violations. In voltage regulation [34, 35] or frequency control [36, 37], a deviation from the desired voltage or frequency will be penalized. When using MARL in the context of home energy management tasks where multiple electrical appliances within a building are controlled by multiple agents, penalties are used to reduce dependence on the main grid [38] or maintain comfort [39].
However, such a global reward induces non-stationarity: All agents adjust their policy simultaneously to obtain a higher reward. Since the reward depends on the decisions of all agents, the individual faces a moving target problem [40]. A common remedy for this problem is to use global observations during training, a paradigm that is referred to as centralized training, decentralized execution (CTDE).
In the context of MARL for coordinated economic dispatch of multiple entities, privacy concerns make it more difficult to share global observations among agents. Some works still use the CTDE paradigm with the justification that training is performed in simulation, where additional information is readily available [41, 42]. An alternative is to only share selected [32] or encoded [43] information between the agents. This requires establishing a communication protocol that can be realized in CommonPower by extension.
II-C Tools for Modeling Power Systems and Interfacing Single-Agent RL
Andes_gym [44] provides a single-agent RL interface for frequency and voltage control. It is based on the ANDES library [45], which features a symbolic modeling framework and optimized numerical simulations. Gym-ANM [46, 47] targets economic dispatch use cases and includes an MPC as a baseline. Grid2op [48, 49] presents a framework for power grid management in which agents can control both the grid topology and power dispatch. Furthermore, it can model opponents that attempt to destabilize the system. PowerGym [50] is a library designed for Volt-Var control in distribution networks and utilizes the Python version of OpenDSS for solving provided system constraints. The tool python-microgrid [51] is a lightweight framework for economic dispatch in microgrids. RLGC [52] is a library tailored for emergency control, e.g., generator dynamic braking or under-voltage load shedding. Lastly, SustainGym [53] is a collection of benchmarks for single-agent and multi-agent control covering five distinct energy management use cases.
II-D Tools for Modeling Power Systems and Interfacing Multi-Agent RL
Toolboxes for MARL exist mainly for home energy management and economic dispatch applications. CityLearn [2, 3] provides both benchmark environments and baseline implementations for demand response of buildings using MARL. This framework is extended with a power grid model in GridLearn [4] so that tasks such as voltage regulation can be addressed. Both libraries focus on decentralized control of active storage components, such as thermal energy storage or batteries, and rely on pre-simulated heating and cooling demands. PowerGridWorld [5] provides a modular framework for modeling arbitrary multi-agent scenarios, including effects of decisions on passive components such as thermal energy storage capacity of buildings.
TABLE I: CommonPower compared to existing Python libraries
Name Focus area
Flexible
component modeling
Symbolic
equations
Decentralized
control
Flexible
prediction models
Andes_gym [44, 45] Stability control
CityLearn [3, 2] Building management
Gym-ANM [47, 46] Economic dispatch
Grid2op [48, 49] Sequential grid operation
GridLearn [4] Building management
PowerGridworld [5] Economic dispatch
PowerGym [50] Voltage control
python-microgrid [51] Economic dispatch
RLGC [52] Emergency control
SustainGym [53] Energy management
CommonPower Economic dispatch
III Preliminaries
III-A Problem Statement
The primary application areas of CommonPower are smart home energy management, demand response, and economic dispatch in microgrids. We model these control tasks as discrete-time receding-horizon optimal control problems. Within a network, we assume a set of nodes $\mathcal{N}$. A node can either be a prosumer or a balancing asset. The set of all prosumers $\mathcal{P}\subseteq\mathcal{N}$ comprises generators, inflexible loads, flexible loads, storage systems, and mixtures thereof. A set of balancing assets $\mathcal{A}=\mathcal{N}\setminus\mathcal{P}$ represents assets that are utilized by the grid operator to establish grid stability. Prosumers can form disjoint coalitions $\mathcal{G}^{i}\subseteq\mathcal{P}$, $\bigcap\mathcal{G}^{i}=\emptyset$, $\bigcup\mathcal{G}^{i}=\mathcal{P}$ in which all members are jointly controlled. We refer to centralized control for the special case in which there are no balancing assets, and all prosumers are controlled centrally, i.e., there is only one coalition $\mathcal{G}=\mathcal{P}=\mathcal{N}$. In every coalition $i$,
  • $u_{t}^{i}$ is the vector of controllable inputs,
  • $\hat{d}_{t}^{i}$ is the vector of uncontrollable exogenous inputs, which is a possibly inaccurate forecast $\forall t>0$,
  • $x_{t}^{i}$ is the state vector,
  • $h^{i}(x_{t}^{i},\hat{d}_{t}^{i})$ describes the input coupling, i.e., constraints arising from power flow or power balance considerations,
  • $f^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i})$ models the system dynamics, and
  • $C^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i})$ is the cost function.
Inputs and states are restricted by constant technical limits $[\underline{u}^{i},\bar{u}^{i}]$ and $[\underline{x}^{i},\bar{x}^{i}]$, respectively. For every time step $t$, each coalition $i$ solves for a control horizon $T$ the optimization problem
$\min_{u^{i}}\sum_{t=0}^{T}C^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i})$ (1a)
$s.t.$
$\begin{split}&u_{t}^{i}\in[\underline{u}^{i},\bar{u}^{i}]\\ &u_{t}^{i}=h^{i}(x_{t}^{i},\hat{d}_{t}^{i})\\ &x_{t}^{i}\in[\underline{x}^{i},\bar{x}^{i}]\\ &x_{t+1}^{i}=f^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i}).\end{split}$ (1b)
In the case of centralized control, all nodes and their input coupling are contained in (LABEL:oc_high_level_coalition_constraints), resulting in a dispatch solution that satisfies all system constraints.
In decentral control, (LABEL:oc_high_level_coalition_constraints) only contains input coupling constraints within coalitions but not between them since we assume that coalitions do not communicate with each other. This means that the independent dispatch decisions of all coalitions might not fulfill all network-level constraints. Therefore, in practice, we always model at least one balancing asset in decentral control. Balancing assets are dispatched as a second stage by solving (1) with a single coalition $\mathcal{G}^{i}=\mathcal{A}$. Here, the previously computed input trajectories for all coalitions are concatenated in $u^{\mathcal{P}}$, which is treated as uncontrollable input. The input coupling constraint is then given by $u_{t}^{\mathcal{A}}=h^{\mathcal{A}}(x_{t}^{\mathcal{A}},\begin{bmatrix}\hat{d}^{\mathcal{A}}&u_{t}^{\mathcal{P}}\end{bmatrix}^{\top})$.
III-B Theoretical Background
The problem statements in Sec. III-A can both for the central and decentral case be considered a sequential decision-making process under uncertainty. RL is the standard machine learning approach for solving such problems. Compared to classic control approaches such as optimal control, RL does not require an explicit definition of the system dynamics $f$. Instead, implicit knowledge about $f$ is gathered through interaction with the system. For the machine learning aspects, we refer to the system as an environment and the RL controller as an agent.
Decentral control of a system with multiple agents can be realized using multi-agent reinforcement learning (MARL). In MARL, the underlying control problem is commonly modeled as a partially observable Markov game (POMG). It is defined as a tuple $(\mathcal{L},\mathcal{S},(\mathcal{U}^{i},\mathcal{O}^{i},R^{i})_{\forall i\in\mathcal{L}},T,\gamma)$ [54], where
  • =$\mathcal{L}=\{1,...,L\}$ is the set of agents,
  • $\mathcal{S}:=[\underline{s},\bar{s}]$ is the global state space of the environment,
  • $\mathcal{O}^{i}:=[\underline{o}^{i},\bar{o}^{i}]$ is the observation space of an agent,
  • $\mathcal{U}^{i}:=[\underline{u}^{i},\bar{u}^{i}]$ is the input space (also referred to as the action space) of an agent,
  • $R^{i}:\mathcal{O}^{i}\times\mathcal{U}^{i}\rightarrow\mathbb{R}$ is the agent-specific reward function,
  • $T:\mathcal{S}\times\mathcal{U}_{1}\times...\times\mathcal{U}_{L}\times\mathcal{S}\rightarrow\mathbb{R}$ is the probability density function modeling state transitions, and
  • $\gamma\in[0,1)$ is the discount factor used to weigh future rewards.
We use the notation $(\mathcal{U}^{i},\mathcal{O}^{i},R^{i})_{\forall i\in\mathcal{L}}$ to refer to the tuple of individual quantities $(\mathcal{U}_{1},...,\mathcal{U}_{L},\mathcal{O}_{1},...,\mathcal{O}_{L},R_{1},...,R_{L})$. Centralized RL control is a special case of the POMG with $L=1$, resulting in a Markov decision process (MDP).
Even though the presented POMG does not model exactly the same problem as (1), the two are closely related. During training, the RL agents learn the policy $\pi^{i}$ with $u^{i}_{t}~{}=~{}\pi^{i}(o^{i}_{t})$ maximizing the expected cumulative discounted reward $\mathbb{E}_{\pi^{i}}[\sum_{k=0}^{T}\gamma^{k}R^{i}_{t+k+1}]$ [55]. However, vanilla RL cannot consider the constraints (LABEL:oc_high_level_coalition_constraints) except for the input limits which are captured by the action space $\mathcal{U}^{i}$. This is why we require a safeguard like the one described in Sec. V-B. Without any safety considerations, the reward is the negative stage cost $R^{i}=-C^{i}$. In Sec. V-B we describe how to modify the reward function to incorporate information about any interventions of the safeguard. We choose the observation of an agent as the concatenation of its state variables and the forecasts of uncontrollable inputs $o^{i}_{t}=\begin{bmatrix}x^{i}_{t}&\hat{d}^{i}_{[t,T]}\end{bmatrix}^{T}$.
III-C Model-based Safeguarding
Among the numerous approaches for safeguarding RL, model-based strategies are particularly well suited for power systems, as they do not require the derivation of auxiliary functions and scale to high dimensions. These methods, often referred to as safety shields or safety filters, restrict the actions of the RL agent to the set of safe inputs at runtime [10, 56, 57]. The set of safe inputs is based on a possibly simplified model of the system and can be computed from a robust control invariant state set [58] or, more flexibly, be formulated as a model predictive control problem [57].
Corresponding strategies can be divided into preemptive and post-posed approaches. Preemptive methods adjust the action selection process of the RL algorithm itself, whereas post-posed methods adjust the action selected by the agent before applying it to the environment. Since our framework is algorithm-agnostic, we only provide an abstraction and corresponding implementation for post-posed approaches, as detailed in Sec. V-B.
IV Modeling
IV-A Power System Entities
CommonPower comprises two domains: the object domain and the symbolic domain. The object domain makes it possible to conveniently compose power systems and provides interfaces to external libraries or tools. In the symbolic domain, all power system entities have a symbolic representation. They specify the variables and constraints that constitute the system model corresponding to the problem formulation in (1). With this approach, object-oriented programming features, such as inheritance, can be leveraged while maintaining full symbolic expressiveness.
IV-A1 Object domain
In CommonPower, physical entities such as buses, electrical devices, or lines are all referred to as power system entities. CommonPower furthermore abstracts the way power flow is modeled as an entity. The implemented entities are described in the following list, and their relationships are visualized in Fig. 2:
  • System: This class is the unique root of the power system representation. It keeps a list of all buses and lines in the system.
  • Bus: This represents a generic bus and can have subordinate child components. By default, a bus comes with an internal power balance equation and sums the costs of its children.
  • Component: A component represents a specific electrical device, e.g., battery, electric vehicle, gas turbine, etc. It does not have children. Implementing sub-classes of components enables flexible modeling of arbitrary devices.
  • Line: Represents a power transmission line connecting two buses. Sub-classes enable flexible modeling in conjunction with corresponding power flow models.
  • Power Flow: This defines how power flow is modeled on the system level, e.g., via power balance, DC power flow constraints, or AC power flow constraints.
Figure 2: UML block definition diagram of power system entities.
IV-A2 Symbolic domain
In the symbolic domain, CommonPower uses the optimization framework Pyomo [59, 60]. The reasoning behind choosing Pyomo over a more generic symbolic library is that the resulting models can readily be used in optimization problems. This is required for optimal control and RL safeguarding. Pyomo uses the concepts of variables Var and parameters Param to differentiate between decision variables and model parameters. To enable a high level of model expressiveness, CommonPower defines a range of model element types that each map to either variables or parameters. Accordingly, these types can be thought of as providing metadata to Pyomo. Additionally, there exist types directly corresponding to constraints and sets in Pyomo. Note that Pyomo can represent arbitrary nonlinear constraints, which makes it possible to model non-convex dynamics or power flow. Specifically, piece-wise linear functions are often useful in power system modeling, e.g., for the approximation of battery dynamics [61]. CommonPower models such occurrences via mixed integer equations and provides a utility that largely automates the generation of corresponding auxiliary constraints (inspired by [62]). This approach generally leads to mixed-integer quadratically constrained programs, which can be solved more efficiently than general nonlinear programs. For illustration, the model of a simple battery storage system is presented in Tab. II. This model introduces a binary charging indicator variable and corresponding constraints based on the big-M method to represent the piece-wise linear dynamics of batteries.
TABLE II: Battery model. $M$ is a large positive constant.
description name type constraint expression/domain
active power $p$ input $p\in\mathbb{R}$
charging indicator $ec$ state $ec\in\{0,1\}$
cost $cost$ state $cost\in\mathbb{R}$
state of charge $soc$ state $soc\in\mathbb{R}_{0}^{+}$
initial soc $soc^{init}$ constant $soc^{init}\in\mathbb{R}_{0}^{+}$
cost of wear $\rho$ constant $\rho\in\mathbb{R}_{0}^{+}$
charge efficiency $\eta^{c}$ constant $\eta^{c}\in[0,1]$
discharge efficiency $\eta^{d}$ constant $\eta^{d}\in[0,1]$
self-discharge $\eta^{s}$ constant $\eta^{s}\in[0,1]$
indicator constraint 1 constraint $\forall t:p_{t}\geq-M(1-ec_{t})$
indicator constraint 2 constraint $\forall t:p_{t}<Mec_{t}$
state initialization constraint $soc_{0}=soc^{init}$
dynamics function constraint $\forall t\in[0,T-1]:soc_{t+1}=\eta^{s}soc_{t}+\eta^{c}(ec_{t})p_{t}+\frac{1}{\eta^{d}}(1-ec_{t})p_{t}$
cost function constraint $\forall t:cost_{t}=\rho(ec_{t})p_{t}-\rho(1-ec_{t})p_{t}$
IV-B Global Symbolic Model
After having modeled a scenario in the object domain, the system is represented by an object tree of entities, with every entity defining its individual symbolic model. Before running a simulation, CommonPower automatically aggregates this distributed symbolic representation into a global symbolic model in the form of (1). To this end, the object tree is traversed and all local symbolic models are added to a hierarchical model in which the original tree structure is maintained via Pyomo blocks. This hierarchical structure can readily be used for optimization within Pyomo and defines the scope of optimal controllers and safeguards in a decentral control scenario.
IV-C Forecasters
CommonPower provides maximum flexibility by defining high-level interfaces for data sources and forecasting algorithms. Forecasters return a prediction for the value of a variable for every time step within the control horizon while having access to past and present values of a set of features. Features can be measurements and additional information, such as the day of the week. In economic dispatch, the quality of forecasts significantly influences the final system cost. To enable an explicit treatment of the associated uncertainty in the safeguarding mechanism, forecasters can return forecast bounds. These bounds signify upper limits $\bar{d}_{t}$ and lower limits $\underline{d}_{t}$, which we treat as guarantees obtained by conformance checking [63]. By default, we assume guaranteed least-conservative bounds, i.e., bounds based on the absolute difference between forecast and the true value, formally $\hat{d}\in[d_{t}+|d_{t}-\hat{d_{t}}|,d_{t}-|d_{t}-\hat{d_{t}}|]$.
CommonPower implements a framework for training, evaluating, and tuning machine learning models to facilitate the study of data-driven forecasting approaches. To obtain a trained forecasting model on a given data source, the user only needs to provide a custom model that implements PyTorch’s nn.Module interface [64] and some parameters for the training pipeline.
V Control
V-A Unified Control Approach
To enable both the deployment and the training of single-agent and multi-agent systems with heterogeneous controllers, CommonPower offers one unified interface based on the gymnasium API [18]. Developing our own interface was necessary because no standard environment representation has emerged for multi-agent RL thus far. We therefore have an internal environment representation and use wrapper classes to adapt to representations used by the respective RL libraries such as StableBaselines [65] or MAPPO [54].
Our interface realizes the interaction of any control algorithm with the underlying power system. The control flow for one episode with $T$ time intervals is provided by Alg. 1: In each time step $t$, the system loops over all controllers, which compute the control inputs for their respective subsystems. During training, the RL agents will sample these control inputs from their current policies. When training has been completed, each RL agent predicts the control inputs deterministically based on its learnt policy. We refer to this mode as deployment. Subsequently, all control inputs are applied simultaneously to advance the discrete-time system dynamics and transition to the next state.
Algorithm 1 Single run
$t\leftarrow t_{0}$
$x_{t}\leftarrow$ init() $\triangleright$ Initialize state and parameters
while $t<t_{0}+T$ do
     $d_{t}\leftarrow$ update_data(t) $\triangleright$ Obtain exogenous inputs
     $\hat{d}_{[t+1,t+T]}\leftarrow$ forecasts(t)
     $o_{t}\leftarrow\begin{bmatrix}x_{t}&&d_{t}&\hat{d}_{[t+1,t+T]}\end{bmatrix}$$\triangleright$ Collect observations
     $u_{t}\leftarrow$ get_controller_inputs($o_{t}$)
     $c_{t}\leftarrow C(x_{t},d_{t},u_{t})$$\triangleright$ Compute cost
     $x_{t+1}\leftarrow f(x_{t},d_{t},u_{t})$$\triangleright$ Execute system dynamics
     $x_{t+1}\leftarrow$ additional_updates(t)
     $t\leftarrow t+1$triangleright Step system forward
end while
We want to demonstrate the effectiveness of our unified approach with a small example. The pseudo-code in Fig. 3 shows how to use CommonPower to set up a power system consisting of two buildings. One is controlled by an RL agent, the other by an optimal controller. To the best of our knowledge, no existing tool is able to realize such a setup with heterogeneous controllers. With CommonPower, the user only has to specify the safeguarding mechanism and the RL algorithm. The RL agent can be trained with a single line of code. After training, the system with the two controllers can be simulated using the same interface.
house_1 = Bus()
house_2 = Bus()
sys = PowerSystem()
sys.add_node(house_1).add_node(house_2)
ctrl_1 = RLController(SafetyShieldX())
ctrl_2 = OptimalController()
# decentralized control structure
ctrl_1.add_entity(house_1)
ctrl_2.add_entity(house_2)
# training
TrainingRunner(sys, RLAlgorithmX()).run()
# deployment
DeploymentRunner(sys).run(T)
Figure 3: Usage example of CommonPower in pseudo-code.
V-B Safeguarding RL-based Controllers
To train and deploy RL-based controllers in safety-critical real-world applications, such as power systems, constraints of the controlled system have to be fulfilled at all times. In the case of decentral control with shared constraints, we adopt a two-stage approach: First, each agent computes its dispatch, and then the grid operator computes the dispatch of balancing assets to fulfill global constraints. This is in line with existing work [38, 39, 32, 42, 41]. Furthermore, it is privacy-preserving since each coalition can compute its own dispatch subject to local constraints based on purely local observations, and the grid operator only needs information about the total energy consumption.
The safeguard of an individual agent solves an optimal control problem that captures the constraints of the subsystem under control over a horizon $T$. The output of the safeguard, the safe action for the current time step $t$, is passed to the environment as shown in Fig. 1. The unique feature of CommonPower is that the constraints of the subsystem are automatically extracted from the symbolic domain of the respective entities as described in Sec. IV-B.
The safeguarding explicitly considers forecast uncertainty, which requires set-based adjustments to the problem formulation from (1). Instead of considering a specific forecast $\hat{d}_{t}^{i}$, we consider the set of forecasted possible realizations $\hat{\mathcal{D}}_{t}^{i}=[\underline{d}_{t}^{i},\overline{d}_{t}^{i}]$. We write an uncertain state as $\mathcal{X}_{t}^{i}$. We can now reformulate the input constraints based on a set of forecasts and an uncertain state:
$u_{t}^{i}=h(x_{t}^{i},\hat{d}_{t}^{i})\in[\underline{u}^{i},\overline{u}^{i}]\Longrightarrow\,u_{t}^{i}\in\mathcal{U}_{t}^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i}),$ (2)
with
$\begin{split}\mathcal{U}_{t}^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i})\subseteq\bigg{\{}&u\in[\underline{u}^{i},\overline{u}^{i}]\,\Big{|}\,\\ &\exists x\in\mathcal{X}_{t}^{i},\exists d\in\hat{\mathcal{D}}_{t}^{i}\colon u=h(x,d)\bigg{\}}.\end{split}$ (3)
For convenience, we will denote the feasible input set as $\mathcal{U}_{t}^{i}$ in the remainder of this section. Note that the specific computation of the feasible input set depends on the input coupling $h(\cdot)$. If $h(\cdot)$ is linear, we can directly compute $\mathcal{U}_{t}^{i}$ with any set representation that is closed under the operations of linear map and Minkowski sum222 Given two sets $\mathcal{A},\mathcal{B}\in\mathbb{R}^{n}$ and a matrix $M\in\mathbb{R}^{m\times n}$, we define the linear map as $M\mathcal{A}=\{Ma\mid a\in\mathcal{A}\}$, and the Minkowski sum as ${\mathcal{A}\oplus\mathcal{B}=\{a+b\mid a\in\mathcal{A},b\in\mathcal{B}\}}$. . If $h(\cdot)$ is nonlinear, we can find convex overapproximations, as demonstrated, for example, in [66].
Depending on the realization of forecasted values, we might have to apply any input from $\mathcal{U}_{t}^{i}$. This means we must consider the entire feasible input set in the system dynamics $f(\cdot)$. To this end, we reformulate the difference equation from (LABEL:oc_high_level_coalition_constraints) into a difference inclusion:
$x_{t+1}^{i}=f^{i}(x_{t}^{i},\hat{d}_{t}^{i},u_{t}^{i})\Longrightarrow\,\mathcal{X}_{t+1}^{i}\subseteq f^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i},\mathcal{U}_{t}^{i}).$ (4)
Again, the specific set-based computation depends on $f(\cdot)$.
Our implementation in CommonPower exclusively uses intervals as set representations to maximize tractability and requires $f(\cdot)$ to be continuous, injective, and monotonous. Under these conditions, all constraints can be efficiently posed as constraints on the intervals’ upper and lower limits. Note that these conditions are usually not restrictive in practice. CommonPower expects instances of Power Flow entities to return $\hat{\mathcal{D}}_{t}^{i}$ as an interval set. All other functionals are given by the user in the non-set-based form of (1) and automatically converted to their set-based equivalent.
Our safeguarding framework is based on the predictive safety filter from [57] and abstracts all post-posed mechanisms. Specifically, the safeguard finds a safe action $\tilde{u}^{i}_{t}$ based on the proposed action $u^{i}_{t}$ of the RL agent by solving
$\begin{split}&\tilde{u}^{i}_{t}=\operatorname*{argmin}_{u^{\prime}_{t}}\phi(u^{\prime}_{t},u^{i}_{t})\\ &s.t.\\ &u_{t}^{i}\in\mathcal{U}_{t}^{i}\\ &\mathcal{X}_{t}^{i}\subseteq[\underline{x}^{i},\bar{x}^{i}]\\ &\mathcal{X}_{t+1}^{i}\subseteq f^{i}(\mathcal{X}_{t}^{i},\hat{\mathcal{D}}_{t}^{i},\mathcal{U}_{t}^{i}).\end{split}$ (5)
where $\phi(u^{\prime}_{t},u^{i}_{t})$ is a cost function defining the type of mechanism. With $\phi(u^{\prime}_{t},u^{i}_{t})=\|u^{\prime}_{t}-u^{i}_{t}\|_{2}$, the approach is commonly called action projection. In contrast, with $\phi(u^{\prime}_{t},u^{i}_{t})=0$, we obtain random actions with the solver initialization as the source of randomness. This is an example of so-called action replacement [14].
When using post-posed safeguarding during RL training, the behavior policy used to gather data differs from the target policy that is being learned. Normally, a policy is updated based on a batch of tuples $(o^{i}_{t},u^{i}_{t},o^{i}_{t+1},r^{i}_{t})$. When a correction of $u^{i}_{t}$ becomes necessary, this tuple could be changed to $(o^{i}_{t},\tilde{u}_{t}^{i},o^{i}_{t+1},r^{i}_{t})$ which features the safe action and the reward obtained from applying this action. However, this would mean updating the policy with actions that do not stem from the most recent policy, which is an expected procedure for off-policy algorithms but can be problematic for on-policy algorithms [14]. Instead, we add an adaption penalty to the reward
$\tilde{r}^{i}_{t}=R^{i}(o^{i}_{t},u^{i}_{t},o^{i}_{t+1})+R^{i}_{\textnormal{penalty}}(u^{i}_{t},\tilde{u}_{t}^{i})$ (6)
such that the tuple used for learning is $(o^{i}_{t},u^{i}_{t},o^{i}_{t+1},\tilde{r}^{i}_{t})$. This penalty informs the agent about the intervention of the safeguard. How to compute the penalty is a design choice and can require careful tuning.
VI Experiments
Figure 4: Mean rollout length during training for four selected combinations of safeguarding approach and penalty. We provide the mean and standard deviation over five random seeds for each configuration. One rollout comprises three consecutive days, meaning that the maximum rollout length is $72$.
To demonstrate the capabilities of CommonPower, we present a benchmark that compares RL agents trained without safeguarding and with different types of safeguards against an optimal control baseline. The scenario is an energy management system for a battery that optimizes the self-consumption of electricity generated by a photovoltaic module. We employ perfect forecasts, which avoids the influence of forecast uncertainty and enables us to highlight differences in performance between the safeguarding approaches. PPO is used as an RL algorithm across all experiments.
VI-A System Model
Figure 5: State of charge of the battery over one week, starting 30 July 2016. We compare three configurations: The optimal controller, an agent trained without safeguard (but deployed with the projection safeguard), and an agent trained with the projection safeguard.
We model a household with an inflexible load, a battery storage system, and a photovoltaic generator. The task of the controller is to optimize the battery usage in order to minimize electricity costs for the household. The buying price for one kWh is set to $0.37$€ and the selling price to $0.08$€. For the household load data as well as the renewable generation, we use time series provided by the SimBench [67] benchmark data set.
VI-B RL Training
In each time step $t$, the RL agent gets an observation
$o_{t}~{}=~{}\begin{bmatrix}soc_{t}&\hat{\ell}_{[t,T]}^{i}&\hat{g}_{[t,T]}&\hat{\varphi}_{[t,T]}\end{bmatrix}^{T}$
where $soc_{t}$ is the state of charge of the battery and ^$\hat{\ell}_{[t,T]}$, $\hat{g}_{[t,T]}$, and $\hat{\varphi}_{[t,T]}$ are the forecasted load, generation, and electricity prices. We choose a forecast horizon of 12 time steps and a time interval of one hour. The action of the agent is the charging/discharging power of the battery, $u_{t}=\begin{bmatrix}p_{t}^{B}\end{bmatrix}.$ The reward $\tilde{r}_{t}$ consists of two components as described in Eq. (6): the negative electricity cost $R_{t}=-p_{t}\,\varphi_{t}$ and a penalty $R_{\textnormal{penalty}}$.
When training without safeguarding, we terminate the episode and reset the system as soon as a constraint – here the maximum or minimum battery capacity – is violated. We evaluate two different types of safeguards:
  • action projection, where the safeguard finds the safe action that is closest to the one proposed by the RL agent, and
  • action replacement, where a random safe action is chosen.
Each safeguarding option is combined with different penalty options. We either add
  • no penalty or
  • a constant penalty of magnitude $R_{\textnormal{penalty}}=-1.0$
to the reward. In addition, we experiment with a distance depending penalty. When using action projection safeguarding, this penalty is proportional to the distance between the safe and the unsafe action, such that $R_{\textnormal{penalty}}=-w\|\tilde{u}_{t}-u_{t}\|_{2},$ where $w$ is a scaling factor. We did not investigate this option for the replacement safeguard as it does not provide any useful information to the agent. For training without safeguarding, the penalty is proportional to the percentage of completed steps before a constraint violation occurs.
TABLE III: Deployment results over 5 training seeds and 5 evaluation periods, each one week long. We report the cumulative cost over the five weeks as a mean over the training seeds for the RL controllers. The mean number of interventions refers to how often the safeguard had to intervene during one evaluation period to prevent constraint violations.
Costs Interventions
Optimal controller 18.55 0.0
RL controller trained without safeguard no penalty 24.58 2.44
constant penalty 22.78 2.96
dd penalty 26.08 2.2
RL controller trained with projection safeguard no penalty 24.50 2.32
constant penalty 26.96 0.08
dd penalty 25.47 3.2
RL controller trained with replacement safeguard no penalty 27.05 0.4
constant penalty 26.57 0.12
dd penalty n/a n/a
VI-C Results and Discussion
Each configuration of RL agents is trained over five random seeds. Fig. 4 visualizes the average rollout length during training for some selected combinations of safeguard and penalty. When training without a safeguard, constraint violations occur frequently and induce a reset of the environment. Therefore, the mean rollout length is shorter. Fig. 4 illustrates that the rollout length can be increased by using a penalty whenever we have to reset.
During deployment, we no longer allow constraint violations and add an action projection safeguard to all RL agents trained without safeguarding. While the state of charge of the battery was randomly initialized during training, it is fixed to half of the capacity during deployment. Tab. III reports the sum of the electricity costs over five evaluation periods, starting February 1, April 15, July 30, September 23, and November 10, and lasting one week. Furthermore, Tab. III displays the mean number of interventions during one evaluation period, with the theoretical maximum being $7*24=168$ interventions. The results are averaged over the five training seeds.
The best performance can be achieved with the RL agents trained without a safeguard and with a constant penalty. Fig. 5 shows that an agent with this configuration has learned a slightly more varying behavior compared to an agent trained with a projection safeguard and a constant penalty. The latter exhibits a very similar charging and discharging behavior for every day, regardless of the actual amount of surplus energy. We hypothesize that the agent trained without a safeguard achieved better exploration because its behavior policy was the same as the policy being optimized during training. This is a central paradigm of on-policy algorithms such as PPO, and adding a safeguard during training changes it. Further experiments are required to establish whether this phenomenon prevails. In any case, a safeguard is necessary whenever interaction with a real system is required during training.
Furthermore, Tab. III shows that using a constant penalty when training with a safeguard can decrease the number of safeguard interventions during deployment. This can induce a rise in costs, possibly because the maximum and minimum capacity of the battery are exploited less. However, obtaining an agent that relies as little as possible on the safeguard can be desirable, for example, when real-time requirements exist and safeguarding involves time-consuming computations.
The experiments demonstrate a complex interaction between the safeguards, the safety penalty, and the ability of the agent to learn cost optimization versus minimization of safety interventions. Investigations along more dimensions, such as further single-agent RL algorithms, multi-agent RL, or imperfect forecasts, are required to enable robust statements on the performance of RL under different safeguarding approaches. With CommonPower, such studies can readily and consistently be conducted across a variety of scenarios.
VII Conclusion
We introduce a comprehensive toolbox, CommonPower, for the study of safe RL controllers in smart grids. CommonPower offers convenient modeling of power systems and serves as a modular framework facilitating comparisons of different safeguarding approaches, levels of decentralization, controller types, and forecasters. Leveraging a symbolic representation of the system, CommonPower automatically extracts system constraints, which serves as the basis for, firstly, a robust model-based safeguarding framework that explicitly considers inaccurate forecasts and, secondly, the automatic derivation of a model predictive controller that can readily serve as a credible baseline for performance evaluation. Moreover, CommonPower enables the simulation of single-agent and multi-agent deployments via a unified interface as well as the combination of RL-based and non-RL controllers in the same system. Planned extensions of CommonPower are the implementation of further entity models, hierarchical control, contingency constraints, further safeguarding approaches, and explicit modeling of energy trading.